Sebastian Kirsch^{– origami art –}^{Marhöferstraße 23A}^{66978 Clausen}^{Germany}  ^{e-mail:} ^{skirsch@t-online.de}^{phone:}  ^{+49 6333 4653}^{fax:}  ^{+49 6333 7222}^{Sebastian Kirsch^{– origami art –}^{Marhöferstraße 23A}^{66978 Clausen}^{Germany}  ^{e-mail:} ^{skirsch@t-online.de}^{phone:}  ^{+49 6333 4653}^{fax:}  ^{+49 6333 7222}}^{Sebastian Kirsch^{– origami art –}^{Marhöferstraße 23A}^{66978 Clausen}^{Germany}  ^{e-mail:} ^{skirsch@t-online.de}^{phone:}  ^{+49 6333 4653}^{fax:}  ^{+49 6333 7222}}^{Sebastian Kirsch^{– origami art –}^{Marhöferstraße 23A}^{66978 Clausen}^{Germany}  ^{e-mail:} ^{skirsch@t-online.de}^{phone:}  ^{+49 6333 4653}^{fax:}  ^{+49 6333 7222}}^{Sebastian Kirsch^{– origami art –}^{Marhöferstraße 23A}^{66978 Clausen}^{Germany}  ^{e-mail:} ^{skirsch@t-online.de}^{phone:}  ^{+49 6333 4653}^{fax:}  ^{+49 6333 7222}}  ^{Sebastian Kirsch^{– origami art –}^{Marhöferstraße 23A}^{66978 Clausen}^{Germany}  ^{e-mail:} ^{skirsch@t-online.de}^{phone:}  ^{+49 6333 4653}^{fax:}  ^{+49 6333 7222}}^{Sebastian Kirsch^{– origami art –}^{Marhöferstraße 23A}^{66978 Clausen}^{Germany}  ^{e-mail:} ^{skirsch@t-online.de}^{phone:}  ^{+49 6333 4653}^{fax:}  ^{+49 6333 7222}}^{Sebastian Kirsch^{– origami art –}^{Marhöferstraße 23A}^{66978 Clausen}^{Germany}  ^{e-mail:} ^{skirsch@t-online.de}^{phone:}  ^{+49 6333 4653}^{fax:}  ^{+49 6333 7222}}^{Sebastian Kirsch^{– origami art –}^{Marhöferstraße 23A}^{66978 Clausen}^{Germany}  ^{e-mail:} ^{skirsch@t-online.de}^{phone:}  ^{+49 6333 4653}^{fax:}  ^{+49 6333 7222}}^{Sebastian Kirsch^{– origami art –}^{Marhöferstraße 23A}^{66978 Clausen}^{Germany}  ^{e-mail:} ^{skirsch@t-online.de}^{phone:}  ^{+49 6333 4653}^{fax:}  ^{+49 6333 7222}}